# Exotic Hidden-heavy Hadrons and Where to Find Them

### v1

@@ Abstract
Physics beyond the Standard Model may include long-lived heavy particles with $\bm{8}$ color charges, such as the gluino in some supersymmetric models. After being created in a collider, a gluino can hadronize before it decays. The ATLAS and CMS collaborations have searched for long-lived gluinos at the Large Hadron Collider [34, 35, 36, 37, 38, 39, 40, 41]. The analyses depend on assumptions about the gluino-hadron spectrum [42, 43, 44], which is the $\bm{8}$-hadron spectrum up to an additive constant associated with the gluino mass. The exotic hidden-heavy hadrons provide constraints on that spectrum. The masses of $X_{c}$ and $Z_{b}$ suggest that the lightest gluino-meson is isospin-0 but isospin-1 is not much heavier. The masses of the $c\bar{c}$ pentaquarks $P_{c\bar{c}}^{+}(4312)$ discovered in 2019 [45] and $P_{c\bar{c}s}^{+}(4338)$ discovered in 2023 [46] suggest that the lightest gluino-baryon is isospin-$\frac{1}{2}$ but isospin-0 is not much heavier. Quantitative B-O analyses of exotic hidden-heavy hadrons that take into account heavy-hadron spin splittings could determine the lowest-energy states in the $\bm{8}$-hadron spectrum and enable more accurate constraints on long-lived gluinos.

### v3

@@ Abstract
Physics beyond the Standard Model may include long-lived heavy particles with $\bm{8}$ color charges, such as the gluino in some supersymmetric models. After being created in a collider, a gluino can hadronize before it decays. The ATLAS and CMS collaborations have searched for long-lived gluinos at the Large Hadron Collider [37, 38, 39, 40, 41, 42, 43, 44]. The analyses depend on assumptions about the gluino-hadron spectrum [45, 46, 47, 48, 49], which is the $\bm{8}$-hadron spectrum up to an additive constant associated with the gluino mass. The exotic hidden-heavy hadrons provide constraints on that spectrum. The masses of $X_{c}$ and $Z_{b}$ suggest that the lightest gluino-meson is isospin-0 but isospin-1 is not much heavier. The masses of the $c\bar{c}$ pentaquarks $P_{c\bar{c}}^{+}(4312)$ discovered in 2019 [50] and $P_{c\bar{c}s}^{+}(4338)$ discovered in 2023 [51] suggest that the lightest gluino-baryon is isospin-$\frac{1}{2}$ but isospin-0 is not much heavier. Quantitative B-O analyses of exotic hidden-heavy hadrons that take into account heavy-hadron spin splittings could determine the lowest-energy states in the $\bm{8}$-hadron spectrum and enable more accurate constraints on long-lived gluinos.